OpenStudy (anonymous):

Beulah took out a loan for $1225 at an 11.4% APR, compounded monthly, to buy a foosball table. If she will make monthly payments of $101.75 to pay off the loan, how many total payments will she have to make?

Miracrown (miracrown):

Ah, is this a business math type of prob? If so, you can simply input the given info into the finance solver section in your calculator. Your Pmt (payments) = 0 as this is what you're trying to find.

Miracrown (miracrown):

You can do it manually too. The calculator method is just eh.. more convenient.

OpenStudy (anonymous):

we can use a formula m= monthly payment P= principal i = monthly interest rate n = number of payments \[ \large m = \frac{Pi(1+i)^{n}}{(1+i)^{n}-1}\] we are given P, m, i . solve for t solving I get n is equal to 12.845 , round up to 13

OpenStudy (anonymous):

as miracrown pointed out, you can use a calculator that has financial capability

OpenStudy (anonymous):

Using a TI 84 calculator, you can go to TVM solver I% = 11.4 PV = 1225 PMT = -101.75 FV = 0 P/Y=12 C/Y=12 move highlighted square to N, and press 2nd Enter for solve N = 12.84589, which agrees with wolfram